Okay, so heres my problem:
evaluate the following:
100^log6.
How can I solve this without a calculator?

Answer by stanbon(75887)

100^log6
------
= (10^2)^log(6)
--------
= 10^(2log(6))
-----
= 10^log(36)
= 36
===========
cheers,
Stan H.